# Prophet: Prompting Large Language Models with Complementary Answer Heuristics for Knowledge-based Visual Question Answering

### v3

@@ 6 Conclusion
In this paper, we present Prophet—a conceptually simple framework which uses LLMs as the knowledge engine for knowledge-based VQA. To better activate the few-shot learning capacity of LLMs, we introduce a novel paradigm to prompt LLMs with two types of complementary answer heuristics. Extensive ablations, comparative experiments, and comprehensive analyses on four diverse knowledge-based VQA datasets show the superiority of Prophet over all existing state-of-the-art methods. Notably, Prophet can be instantiated with varied combinations of a wide range of VQA models and LLMs, showing its flexibility, scalability, and generalizability. We hope that our work can inspire future research on knowledge-based VQA and universal multimodal learning in the era of LLMs.
Appendix A More Implementation Details
A.1 The Default VQA Model
@@
TABLE VIII: Ablations for training strategies. All variants use the improved model architecture in the last row in Table VII.
@@
TABLE IX: An exemplar prompt for the standard Prophet. We show one in-context example here due to space limitations. Following the implementations in PICa [15] and KAT [16], we use a special symbol ‘$=$’ to separate each two lines.
@@
TABLE X: An exemplar prompt for the Prophet variant on the MC task of A-OKVQA. Compared to the standard prompt in Table IX, we add one extra line of choices for the example and testing input, and change the output format to adapt to the multiple-choice task. All the differences are marked in red.
@@
TABLE XI: An exemplar prompt for the Prophet variant on ScienceQA (IMG). The sentences marked in red are the optional text hints provided by the dataset.
@@
TABLE XII: An exemplar prompt for the Prophet variant on TextVQA. Compared to the standard prompt, we additionally introduce the OCR tokens (marked in red) extracted from an off-the-shelf OCR system.
Training recipe. We first pretrain the model on the augmented train+val+vg dataset from VQAv2 [47] and Visual Genome[52], with excluding the samples whose images are used in the testing split of OK-VQA to avoid data contamination. The settings for the pretraining stage are identical to the original implementation of MCAN. After that, the model is finetuned on the downstream OK-VQA and A-OKVQA datasets, respectively. For finetuning, the commonly used strategy is to replace the last linear layer (i.e., the classification layer) with a new layer to adapt to the answer vocabulary of the downstream dataset. However, the answer vocabularies of the pretraining and finetuning datasets are partially overlapped. To maximally utilize the pretrained model parameters in the last layer, we inherit the parameters of existing answers and append new parameters for the new answers. After that, we freeze all the pretrained parameters and only update the new parameters for one epoch as a warm-up, and then train all model parameters for the rest training epochs.
Table VIII shows the effects of different training strategies. Even without finetuning, the pretrained model (b) is superior to the model trained from scratch (a), implying the importance of pretraining. Moreover, our new finetuning strategy (d) leads to significantly better performance than the commonly used strategy (c), showing the effectiveness of inheriting model parameters for existing answers.
A.2 Prompt Formats
We show an exemplar prompt for the standard Prophet in Table IX and an exemplar prompt for the variant designed for the MC task of A-OKVQA in Table X. The exemplar prompts for ScienceQA and TextVQA are illustrated in Table XI and XII, respectively.
Appendix B More Qualitative and Quantitative Analyses
@@
TABLE XIII: Per-category accuracies of MCAN (stage-1) and Prophet (stage-2). This performance improvements of using GPT-3 are observed on all categories.
We show the per-type accuracies of MCAN (stage-1) and Prophet (stage-2) in Table XIII. Prophet outperforms MCAN on all categories, indicating that generality of the knowledge in GPT-3. The improvement on the “Science and Technology” category is not as large as the rest categories. which can be explained that the required knowledge for this category is more specialized and professional. These questions are also challenging for humans.
@@
TABLE XIV: The distribution of failure causes by human studies.
@@
We perform human studies to analyze the causes of wrong predictions in Table XIV. For each category, we randomly sample 10% testing samples that Prophet fails to get the correct answer. This results in 172 samples. We ask three annotators to categorize each sample into one of the following four failure causes: (a) insufficient visual understanding; (b) incorrect knowledge reasoning; (c) correct but differently expressed answer; (d) others (e.g., the failure is caused by the ambiguity of the question). From the results, we can see that the cause of “(b) incorrect knowledge reasoning” accounts for the highest proportion, which suggests that the bottleneck of Prophet still lies in the knowledge acquisition and reasoning. The cause of “(a) insufficient visual understanding” has the second highest proportion, showing the potential of devising more powerful VQA models. The cause of “(c) correct but differently expressed answer” also accounts for a considerable proportion. This reflects the limitation of the annotations and evaluation metric of OK-VQA.
Figure 6 demonstrates some testing samples from different knowledge categories. In the 1st-3rd columns, we show the correctly answered samples with different prediction behaviors (i.e., keep top-1, in top 2-$K$, and beyond top-$K$). The visualized results indicate that Prophet can adaptively choose suitable answers from candidates. In the last column, we show some failure samples, implying that there is still room for future improvement.

### v4

@@ 5 Conclusion
In this paper, we present Prophet—a conceptually simple framework which uses LLMs as the knowledge engine for knowledge-based VQA. To better activate the few-shot learning capacity of LLMs, we introduce a novel paradigm to prompt LLMs with complementary answer heuristics. Extensive ablations, comparative experiments, and comprehensive analyses on four diverse knowledge-based VQA datasets show the superiority of Prophet over all existing state-of-the-art methods. Notably, Prophet can be integrated with varied combinations of VQA models, LLMs, and even LMMs, showing the flexibility, scalability, and generalizability of our framework.
Appendix A Broader Impact
From the larger multimodal model (LMM) point of view, Prophet is a loosely-coupled LMM consisting of a vision-language (VL) model and a frozen LLM, aiming to endow the VL model with knowledge reasoning ability. Compared with the tightly-coupled LMMs (e.g., Flamingo [11] and LLaVa [42]) which jointly optimize the VL model and LLM in an end-to-end manner, Prophet is more flexible that can support any open-source or commercial LLM.
Moreover, Prophet can also be regarded as a learning-to-prompt paradigm that learns an external model to generate prompts to better comprehend the target task, thus facilitating the capability of the pretrained LLM (or LMM). From this point of view, recent studies like VoxPoser [60] and SoM-Prompting [61] share a similar idea with our work. We believe this paradigm can be widely used in a variety of LLM-related tasks.
Appendix B More Implementation Details
B.1 The Default VQA Model
@@
config setting
optimizer AdamW
weight decay 0.01
optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.98$
batch size 64
warm-up learning rate 1e-3
warm-up strategy only update new parameters
warm-up epochs 1
base learning rate 5e-5
learning rate schedule step decay
learning rate decay rate 0.2
learning rate decay epoch 6
total training epochs 6
TABLE VIII: Training settings. These hyper-parameters are used in our experiments.
@@
TABLE IX: Ablations for training strategies. All variants use the improved model architecture in the last row in Table VII.
@@
TABLE X: An exemplar prompt for the standard Prophet. We show one in-context example here due to space limitations. Following the implementations in PICa [21] and KAT [22], we use a special symbol ‘$=$’ to separate each two lines.
@@
TABLE XI: An exemplar prompt for the Prophet variant on the MC task of A-OKVQA. Compared to the standard prompt in Table X, we add one extra line of choices for the example and testing input, and change the output format to adapt to the multiple-choice task. All the differences are marked in red.
@@
TABLE XII: An exemplar prompt for the Prophet variant on ScienceQA (IMG). The sentences marked in red are the optional text hints provided by the dataset.
@@
TABLE XIII: An exemplar prompt for the Prophet variant on TextVQA. Compared to the standard prompt, we additionally introduce the OCR tokens (marked in red) extracted from an off-the-shelf OCR system.
Training recipe. We first pretrain the model on the augmented train+val+vg dataset from VQAv2 [47] and Visual Genome[48], with excluding the samples whose images are used in the testing split of OK-VQA to avoid data contamination. The settings for the pretraining stage are identical to the original implementation of MCAN. After that, the model is finetuned on the downstream OK-VQA and A-OKVQA datasets, respectively. For finetuning, the commonly used strategy is to replace the last linear layer (i.e., the classification layer) with a new layer to adapt to the answer vocabulary of the downstream dataset. However, the answer vocabularies of the pretraining and finetuning datasets are partially overlapped. To maximally utilize the pretrained model parameters in the last layer, we inherit the parameters of existing answers and append new parameters for the new answers. After that, we freeze all the pretrained parameters and only update the new parameters for one epoch as a warm-up, and then train all model parameters for the rest training epochs. The detailed settings for the finetuning stage are shown in Table VIII.
Table IX shows the effects of different training strategies. Even without finetuning, the pretrained model (b) is superior to the model trained from scratch (a), implying the importance of pretraining. Moreover, our new finetuning strategy (d) leads to significantly better performance than the commonly used strategy (c), showing the effectiveness of inheriting model parameters for existing answers.
B.2 Prompt Formats
We show an exemplar prompt for the standard Prophet in Table X and an exemplar prompt for the variant designed for the MC task of A-OKVQA in Table XI. The exemplar prompts for ScienceQA and TextVQA are illustrated in Table XII and XIII, respectively.
Appendix C More Quantitative and Qualitative Analyses
@@
TABLE XIV: Per-category accuracies of MCAN (stage-1) and Prophet (stage-2). This performance improvements of using GPT-3 are observed on all categories.
Stage 1 pred. Stage 2 pred.     correct     wrong
correct 54.4% 4.2%
wrong 12.0% 29.4%
TABLE XV: Prophet’s combinatorial prediction behaviors in two stages. Prophet maintains the majority of correct predictions at stage-1, and the accuracy improvement by stage-2 is mainly because the number of wrong-to-correct samples is larger than that of the correct-to-wrong samples.
C.1 Quantitative Analysis
First, we show the per-type accuracies of MCAN (stage-1) and Prophet (stage-2) in Table XIV. Prophet outperforms MCAN on all categories, indicating that generality of the knowledge in GPT-3. The improvement on the “Science and Technology” category is not as large as the rest categories. which can be explained that the required knowledge for this category is more specialized and professional. These questions are also challenging for humans.
Second, we calculate the distribution of four situations of the predictions from stage-1 and stage-2 in Table XV. From the results, we can see that: (i) Prophet maintains the majority of correct predictions by MCAN and only 4.2% samples are overturned; (ii) the improvement of Prophet is mainly due to the fact that the proportion of wrong-to-correct samples (12.4%) is larger than that of the correct-to-wrong samples (4.2%); (iii) there are still a non-negligible amount of samples (29.4%) that both MCAN and Prophet fail, which leaves sufficient room for future improvement.
@@ C.1 Quantitative Analysis
TABLE XVI: The distribution of failure causes by human studies.
Third, we perform human studies to analyze the causes of wrong predictions in Table XVI. For each category, we randomly sample 10% testing samples that Prophet fails to get the correct answer. This results in 172 samples. We ask three annotators to categorize each sample into one of the following four failure causes: (a) insufficient visual understanding; (b) incorrect knowledge reasoning; (c) correct but differently expressed answer; (d) others (e.g., the failure is caused by the ambiguity of the question). From the results, we can see that the cause of “(b) incorrect knowledge reasoning” accounts for the highest proportion, which suggests that the bottleneck of Prophet still lies in the knowledge acquisition and reasoning. The cause of “(a) insufficient visual understanding” has the second highest proportion, showing the potential of devising more powerful VQA models. The cause of “(c) correct but differently expressed answer” also accounts for a considerable ratio. This reflects the limitation of the annotations and evaluation metric of OK-VQA.
method stage-1 stage-2 accu.
VQA model LMM LLM/ LMM pure text/ img+text
GPT-4o - - - - 46.0
L1.5-7B [63] - - - - 58.3
Q2.5-7B [64] - - - - 59.7
Prophet MCAN - GPT-4o pure-text 62.9
L1.5-7B - GPT-4o pure-text 66.5
Q2.5-7B - GPT-4o pure-text 66.4
Prophet++ MCAN GPT-4o GPT-4o img+text 64.2
L1.5-7B GPT-4o GPT-4o img+text 68.3
Q2.5-7B GPT-4o GPT-4o img+text 67.2
TABLE XVII: More ablation study of the latest LMMs in Prophet and Prophet++. The first split contains the direct-prediction results of three state-of-the-art LMMs, namely GPT-4o, LLaVA-1.5-7B (L1.5-7B) [63], and Qwen2.5-VL-7B (Q2.5-7B)[64]. The second and last splits contain different combinations of LMMs for Prophet and Prophet++, respectively. $N$=16 is used for Prophet and Prophet++.
Finally, we investigate the combinations of the latest LMMs for Prophet and Prophet++ in Table XVII, which complements the results of Table 3 in the main text. The following observations are obtained from the results: (i) the direct-prediction performance by LLaVA-1.5-7B (L1.5-7B) [63] and Qwen2.5-VL-7B (Q2.5-7B) [64] significantly exceed that of GPT-4o, which may be explained by the fact that OK-VQA has been included in their training data; (ii) by utilizing the strong LMM L1.5-7B (or Q2.5-7B) as the VQA model, the resulting Prophet and Prophet++ significantly outperform their direct-prediction counterparts and MCAN-based counterparts, respectively, which verifies the effectiveness and generalizability of our frameworks to adapt to the latest LMMs; (iii) although Q2.5-7B achieves slightly better direct-prediction accuracy than L1.5-7B (59.7 vs. 58.3), its corresponding Prophet and Prophet++ models are inferior to the counterparts with L1.5-7B. This could be explained by the observation that the answer candidates generated by Q2.5-7B are more likely to be synonyms with similar confidence scores, which may mislead the LLM/LMM in stage-2 and limit the final performance.
@@
C.2 Qualitative Analysis
Fig. 7 demonstrates some testing samples from different knowledge categories. In the 1st-3rd columns, we show the correctly answered samples with different prediction behaviors (i.e., keep top-1, in top 2-$K$, and beyond top-$K$). The visualized results indicate that Prophet can adaptively choose suitable answers from candidates. In the last column, we show some failure samples, implying that there is still room for future improvement.